# Radio observations point to a moderately relativistic outflow in the fast X-ray transient EP241021a

### v1

@@ Abstract
In this paper, we present long-term radio observations of the gamma-ray dark EP241021a, which exhibits sustained radio emission for over 100 days, placing it among the longest-lived radio afterglows. We detect signature of interstellar scintillation in early epochs, allowing us to constrain the angular size and Lorentz factor of the emitting region. Our observations point to an outflow that is at least mildly relativistic with $\Gamma\gtrsim 4$. Afterglow modeling favors a moderately relativistic ($\Gamma\approx$40) and collimated ($\theta_{j}\approx$0.16 rad) outflow interacting with a low-density interstellar medium. The derived beaming-corrected kinetic energy, $E_{\mathrm{k}}\sim 1.6\times 10^{51}$ erg, and low radiative efficiency, $\eta_{\gamma}\lesssim$1%, are consistent with a standard relativistic explosion which did not produce bright gamma-rays. Alternatively, a highly-relativistic ($\Gamma\gtrsim$100) structured jet remains consistent with our observations if seen substantially off-axis ($\theta_{\rm obs}>\theta_{c}$). In the latter case, the initial X-ray flare detected by EP would be caused by the lower-$\Gamma$ ejecta from the lateral wings intercepting our line of sight rather than by traditional prompt-emission mechanisms within the jet core.
keywords:
@@
pubyear: 2025pagerange: Radio observations point to a moderately relativistic outflow in the fast X-ray transient EP241021aA
@@ 1 Introduction
Fast X-ray transients (FXRTs) are sudden and intense bursts of X-ray emission occurring on timescales ranging from a few seconds to several hours (Arefiev et al., 2003; Bauer et al., 2017; Quirola-Vásquez et al., 2022). While FXRTs have been extensively investigated within the Galactic environment (Connors, 1988; Smith et al., 2006; Sguera et al., 2016; Ferrigno et al., 2019), their extragalactic counterparts remain poorly understood.
@@
Extragalactic FXRTs typically exhibit peak luminosities spanning $10^{42}$ to $10^{50}$ erg s-1, placing them among the most luminous transient phenomena in the X-ray regime. Space-based X-ray observatories such as Swift (Burrows et al., 2005), Chandra (Weisskopf et al., 2002), XMM-Newton (Jansen et al., 2001), and eROSITA (Predehl et al., 2021) have played a critical role in their discovery and characterization (e.g., Glennie et al., 2015; Xue et al., 2019; Lin et al., 2020; Ai & Zhang, 2021; Lin et al., 2022). However, the long delay between the transient onset and its identification in archival searches often prevented placing meaningful constraints on their physical origin. A variety of progenitor scenarios, including compact object mergers, supernova shock break-outs, proto-magnetars, or tidal disruption events (TDEs) remain possible (Xue et al., 2019; Novara et al., 2020; Lin et al., 2022; Castaneda Jaimes et al., 2024). A notable case of a FXRT was reported by Soderberg et al. (2008), who discovered the X-ray transient, XRF080109, in NGC 2770 at a distance of 27 Mpc. With a peak luminosity of  $\approx\,6\times 10^{43}\,\mathrm{erg\,s^{-1}}$, XRF080109 featured a 600-s X-ray flash followed by the emergence of supernova SN2008D, providing the first robust link between FXRTs and the core-collapse of massive stars (Soderberg et al., 2008; Modjaz et al., 2008; Mazzali et al., 2008). More recently, (Bauer et al., 2017) reported the discovery of the FXRT CDF-S XT1 in the Chandra Deep Field-South, characterized by a rapid rise of  $\sim\,100\,\mathrm{s}$, well described by a power-law decay with a slope of $\alpha\approx-1.53$, and a relatively hard spectrum with a photon index of 1.43. FXRT CDF-S XT1 was localized to a faint dwarf host galaxy at a redshift of $z=2.76$ (Quirola-Vásquez et al., 2024), suggesting a much more energetic explosion than XRF080109.
@@
The advent of the Einstein Probe (EP) has significantly advanced our understanding of FXRTs. By enabling the real-time discovery and localization of these events, EP has facilitated the identification of dozens of new FXRTs. Some of them, such as EP240219A (Yin et al., 2024a), EP240315A (Liu et al., 2025a; Ricci et al., 2025; Gillanders et al., 2024), and EP240801a (Jiang et al., 2025a), have been linked to classical gamma-ray bursts (GRBs; Kouveliotou et al., 1993; Woosley & Bloom, 2006). However, most FXRTs do not feature a gamma-ray counterpart, i.e., gamma-ray-dark FXRTs (e.g., EP240414a; Sun et al., 2024a). A subset of these gamma-ray dark FXRTs may originate from extremely distant GRBs, whose spectral peaks are redshifted into the X-ray band (Heise, 2003; D’Alessio et al., 2006); in other cases, such as EP240414A ($z=0.4$; Sun et al., 2024a; Srivastav et al., 2025; van Dalen et al., 2024), the events occur at relatively low redshifts yet still lack detectable gamma-ray counterparts, further challenging our understanding of their physical origins. A leading class of models attributes such gamma-ray dark FXRTs to relativistic outflows powered by a central engine. For example, off-axis jets, dirty fireballs, and choked jets associated to either GRBs or TDEs can reproduce the observed timescales and luminosities of FXRTs (Rhoads, 2003; Paczyński, 1998; Dermer, 1999; Mészáros & Waxman, 2001). Alternatively, models involving supernova shock breakouts, classical TDEs, or fast blue optical transients (FBOTs) may produce X-ray transients without invoking relativistic jets (Soderberg et al., 2008; Holoien et al., 2016; Prentice et al., 2018).
@@
In this context, radio observations serve as a powerful diagnostic for distinguishing among these scenarios. Unlike optical emission, the radio band probes fast-moving ejecta from early to very late ($\gtrsim 1$ month) timescales, tracing long-lived synchrotron emission from relativistic electrons in the blast wave. These measurements provide valuable constraints on the energy budget and expansion dynamics of the outflow over extended periods (see e.g., Ricci et al., 2025; Bright et al., 2025). As radio waves propagate through the interstellar medium (ISM), they are subject to various propagation effects, including interstellar scintillation (see e.g., Lee & Jokipii, 1975; Armstrong, 1984).
@@
This phenomenon has been extensively studied in compact radio sources such as pulsars and active galactic nuclei (AGN; Rickett, 1990; Goodman, 1997), and has also been observed in several GRB afterglows, (Frail et al., 1997, 2000; Chandra et al., 2008; Granot, 2008). ISS arises from small-scale electron density fluctuations in the ISM, which may induce rapid variations in signal intensity if the emitting source is sufficiently compact. Therefore, ISS provide important constraints on the angular size of the emitting region and, consequently, on the Lorentz factor of the outflow (Heeschen & Rickett, 1987; Dennett-Thorpe & de Bruyn, 2002; Lovell et al., 2008; Rickett et al., 1995), helping to determine whether the emission arises from a relativistic outflow.
@@
We identify evidence of ISS, offering a propagation-based explanation for the variability seen in the early-time radio light curve, which allows us to probe the physical properties of its relativistic outflow. Combined with constraints from afterglow modeling, these findings strongly support the presence of an outflow that is at least mildly relativistic with Lorentz factor of $\approx$4. We first examine in Section 2 the prompt emission properties. In Section 3, we provide a comprehensive overview of our radio monitoring campaign using ATCA and e-MERLIN. In Section 4 we use the observed scintillation to constrain the source’s angular size and the Lorentz factor. Section 5 presents the results of our afterglow modeling and Section 6 discusses their implications for the physical parameters of the event. Finally, in Section 7, we summarize our results. Throughout this work, we adopt a flat $\Lambda$CDM cosmology model with the matter density parameter $\Omega_{m}=0.315$, and the Hubble-Lemaître constant $H_{0}=67.4~{}\text{km\,s}^{-1}~{}\text{Mpc}^{-1}$ (Planck Collaboration et al., 2020).
@@ 2 Prompt Emission
Our dataset comprises 84 EP events observed up to March 21, 2025. Of these, 55 events have complete prompt emission coverage from the Fermi Gamma-ray Burst Monitor (GBM; Meegan et al., 2009) or were detections from Konus-Wind (Aptekar et al., 1995).
@@
Fig. 2 compares the (observer frame) gamma-ray (10–1000 keV) and X-ray (0.5–4 keV) fluxes for a sample of FXRTs, including EP241021a. For this study, $3\sigma$ upper limits on gamma-ray fluxes were derived using the Fermi/GBM targeted search, an internal tool developed by the GBM team for gravitational-wave follow-up (Blackburn et al., 2015). We adopt a Band function [^1] The Band function is a phenomenological model of GRB spectra, defined as: $N(E)=\begin{cases}A\left(\frac{E}{100\,\mathrm{keV}}\right)^{\alpha}\exp\left(-\frac{E}{E_{0}}\right),&E<(\alpha-\beta)E_{0}\\ A\left(\frac{(\alpha-\beta)E_{0}}{100\,\mathrm{keV}}\right)^{\alpha-\beta}\exp(\beta-\alpha)\left(\frac{E}{100\,\mathrm{keV}}\right)^{\beta},&E\geq(\alpha-\beta)E_{0}\end{cases}$ with the peak energy given by $E_{\rm peak}=(2+\alpha)E_{0}$ (Band et al., 1993). with typical GRB parameters $\alpha=-1.0$, $\beta=-2.3$ and $E_{\rm peak}=230~{}\mathrm{keV}$ (Normal template) to calculate the upper limits of energy fluxes in the case of non-detections. For events with detections, the time-averaged gamma-ray fluxes were taken directly from the values reported through GCN circulars (see Appendix A). The dashed lines in Fig. 2 indicate constant photon indices ($\Gamma=0.5$, $1.0$, $1.5$), assuming a simple power-law spectrum. We observe that most gamma-ray bright FXRTs lie at $\Gamma\sim 1.0$, consistent with the region typically occupied by standard GRBs (Poolakkil et al., 2021).
For EP241021a, we used the X-ray flux (0.5–4 keV) estimated by Hu et al. (2024b) of $(3.3^{+1.5}_{-1.7})\times 10^{-10}~{}\mathrm{erg~{}cm^{-2}~{}s^{-1}}$. Using the Fermi/GBM targeted search, we place a $3\sigma$ upper limit on the gamma-ray flux of EP241021a as $<3.7\times 10^{-8}~{}\mathrm{erg~{}cm^{-2}~{}s^{-1}}$ for a timescale of 8.192 s. At a redshift of $z=0.75$, this corresponds to a limit on the isotropic-equivalent gamma-ray energy of $E_{\gamma,\rm iso}\lesssim 8\times 10^{50}$ erg, consistent with the value inferred by Busmann et al. (2025) and at the lower end of the distribution for cosmological GRBs, although still consistent with it.
@@ 3.1 Australia Telescope Compact Array
Australia Telescope Compact Array (ATCA) observations were conducted over 11 epochs between October 28, 2024, and January 25, 2025, covering frequencies from 2.1 GHz to 16.7 GHz. The bandwidth in all the observations was 2 GHz in the Compact Array Broadband Backend 1 MHz per frequency channel mode. The primary and bandpass calibrator was 1934-638 for all observations at 2.1, 5.5 and 9 GHz. At 16.7 GHz, 1921–293 served as the bandpass calibrator observations, while 1934-638 remained the primary calibrator. The phase calibrator 0146+056 was used to correct complex gains. The 21.2 GHz data were taken at the same time of the 16.7 GHz ones but the 21.2 GHz flux densities are not presented in this work because high system temperature values and large flux density scale corrections from the primary calibrator resulted in unreliable flux density values for the phase calibrator and the target source. Data reduction was performed using the Miriad (Sault et al., 1995) software package, which was used to flag Radio Frequency Interferences (RFIs), calibrate the complex gains, bootstrap the flux density scale, and image the fringe visibilities. A null robustness parameter value was used as a good trade-off between uniform and natural weighting in the imaging strategy. Final cleaned radio maps were inspected by Karma kvis (KARMA kvis, 1995) and ds9 (Smithsonian Astrophysical Observatory, 2000) visualisation softwares. When a detection was found the peak flux was extracted, otherwise, a 3-$\sigma$ upper limit was derived from the Root Mean Square (RMS) noise in the restored maps. The uncertainties on the flux densities were estimated by summing in quadrature the map RMS noise and 5% of the peak flux density, accounting for residual complex gain calibration errors.
@@
EP241021a was observed with enhanced Multi Element Remotely Linked Interferometer Network (e-MERLIN) in four epochs at the centre frequency of 5 GHz with a total bandwidth of 512 MHz divided in 4 sub-bands of 128 MHz each. The data were correlated by the e-MERLIN software correlator. The resulting fringe visibilities were calibrated in the CASA software v.5.8 using the e-MERLIN CASA pipeline (e-MERLIN pipeline, 2019; Moldon, 2021) and imaged in CASA v.5.5.0 using a Briggs weighting scheme and a robustness value of 0.5. The target was observed together with 1331+3030 (alias 3C286), 1407+2827, 0154+0823 as the flux calibrator , the pointing calibrator and the bandpass calibrator, respectively. The target was fit with a 2D Gaussian in CASA and the peak flux was quoted in the flux density measurement, being it point-like in the $100\times 40$ mas final cleaned maps. The e-MERLIN position obtained by 2D Gaussian fitting to the cleaned map of EP241021a is RA = 01h 55m 23.4325s, Dec = +05 56 17.774′′ (J2000) with a positional $1\sigma$ uncertainity of (3, 9) mas in RA and DEC, respectively. Flux uncertainties were estimated as detailed in Table 1. Detections and the earlier flux upper limit obtained by this work are shown in Table 2.

### v2

@@ Abstract
In this paper, we present long-term radio observations of the gamma-ray dark EP241021a, which exhibits sustained radio emission for over 100 days, placing it among the longest-lived radio afterglows. We detect signature of interstellar scintillation in early epochs, allowing us to constrain the angular size and Lorentz factor of the emitting region. Our observations point to an outflow that is at least mildly relativistic with $\Gamma\gtrsim 4$. Afterglow modeling favors a moderately relativistic ($\Gamma\approx$40) top-hat jet ($\theta_{c}\approx$0.16 rad) interacting with a low-density interstellar medium. The derived beaming-corrected kinetic energy, $E_{\mathrm{k}}\sim 1.6\times 10^{51}$ erg, and low radiative efficiency, $\eta_{\gamma}\lesssim$1%, are consistent with a standard relativistic explosion which did not produce bright gamma-rays. Alternatively, a highly-relativistic ($\Gamma\gtrsim$100) structured jet remains consistent with our observations if seen substantially off-axis ($\theta_{\rm obs}>\theta_{c}$). In the latter case, the initial X-ray flare detected by EP would be caused by the lower-$\Gamma$ ejecta from the lateral wings intercepting our line of sight rather than by traditional prompt-emission mechanisms within the jet core.
@@ 1 Introduction
Fast X-ray transients (FXRTs) are sudden and intense bursts of X-ray emission occurring on timescales ranging from a few seconds to several hours (V. A. Arefiev et al., 2003; F. E. Bauer et al., 2017; J. Quirola-Vásquez et al., 2022). While FXRTs have been extensively investigated within the Galactic environment (A. Connors, 1988; D. M. Smith et al., 2006; V. Sguera et al., 2016; C. Ferrigno et al., 2019), their extragalactic counterparts remain poorly understood.
@@
Extragalactic FXRTs typically exhibit peak luminosities spanning $10^{42}$ to $10^{50}$ erg s-1, placing them among the most luminous transient phenomena in the X-ray regime. Space-based X-ray observatories such as Swift (D. N. Burrows et al., 2005), Chandra (M. C. Weisskopf et al., 2002), XMM-Newton (F. Jansen et al., 2001), and eROSITA (P. Predehl et al., 2021) have played a critical role in their discovery and characterization (e.g., A. Glennie et al., 2015; Y. Q. Xue et al., 2019; D. Lin et al., 2020; S. Ai & B. Zhang, 2021; D. Lin et al., 2022). However, the long delay between the transient onset and its identification in archival searches often prevented placing meaningful constraints on their physical origin. A variety of progenitor scenarios, including compact object mergers, supernova shock break-outs, proto-magnetars, or tidal disruption events (TDEs) remain possible (Y. Q. Xue et al., 2019; G. Novara et al., 2020; D. Lin et al., 2022; J. Castaneda Jaimes et al., 2024). A notable case of a FXRT was reported by A. M. Soderberg et al. (2008), who discovered the X‐ray transient, XRF080109, in NGC 2770 at a distance of 27 Mpc. With a peak luminosity of  $\approx\,6\times 10^{43}\,\mathrm{erg\,s^{-1}}$, XRF080109 featured a 600-s X-ray flash followed by the emergence of supernova SN2008D, providing the first robust link between FXRTs and the core-collapse of massive stars (A. M. Soderberg et al., 2008; M. Modjaz et al., 2008; P. A. Mazzali et al., 2008). More recently, (F. E. Bauer et al., 2017) reported the discovery of the FXRT CDF-S XT1 in the Chandra Deep Field-South, characterized by a rapid rise of  $\sim\,100\,\mathrm{s}$, well described by a power-law decay with a slope of $\alpha\approx-1.53$, and a relatively hard spectrum with a photon index of 1.43. FXRT CDF-S XT1 was localized to a faint dwarf host galaxy at a redshift of $z=2.76$ (J. Quirola-Vásquez et al., 2024), suggesting a much more energetic explosion than XRF080109.
@@
The advent of the Einstein Probe (EP; W. Yuan et al. 2022, 2025) has significantly advanced our understanding of FXRTs. By enabling the real-time discovery and localization of these events, EP has facilitated the identification of dozens of new FXRTs. Some of them, such as EP240219A (Y.-H. I. Yin et al., 2024), EP240315A (Y. Liu et al., 2025; R. Ricci et al., 2025; J. H. Gillanders et al., 2024), and EP240801a (S.-Q. Jiang et al., 2025), have been linked to classical gamma-ray bursts (GRBs; C. Kouveliotou et al., 1993; S. E. Woosley & J. S. Bloom, 2006). However, most FXRTs do not feature a gamma-ray counterpart, i.e., gamma-ray-dark FXRTs (e.g., EP240414a; H. Sun et al., 2024a). A subset of these gamma-ray dark FXRTs may originate from extremely distant GRBs, whose spectral peaks are redshifted into the X-ray band (J. Heise, 2003; V. D’Alessio et al., 2006); in other cases, such as EP240414A ($z=0.4$; H. Sun et al., 2024a; S. Srivastav et al., 2025; J. N. D. van Dalen et al., 2024), the events occur at relatively low redshifts yet still lack detectable gamma-ray counterparts, further challenging our understanding of their physical origins. A leading class of models attributes such gamma-ray dark FXRTs to relativistic outflows powered by a central engine. For example, off-axis jets, dirty fireballs, and choked jets associated to either GRBs or TDEs can reproduce the observed timescales and luminosities of FXRTs (J. E. Rhoads, 2003; B. Paczyński, 1998; C. D. Dermer, 1999; P. Mészáros & E. Waxman, 2001). Alternatively, models involving supernova shock breakouts, classical TDEs, or fast blue optical transients (FBOTs) may produce X-ray transients without invoking relativistic jets (A. M. Soderberg et al., 2008; T. W. S. Holoien et al., 2016; S. J. Prentice et al., 2018).
@@
In this context, radio observations serve as a powerful diagnostic for distinguishing among these scenarios. Unlike optical emission, the radio band probes fast-moving ejecta from early to very late ($\gtrsim 1$ month) timescales, tracing long-lived synchrotron emission from relativistic electrons in the blast wave. These measurements provide valuable constraints on the energy budget and expansion dynamics of the outflow over extended periods (see e.g., R. Ricci et al., 2025; J. S. Bright et al., 2025). As radio waves propagate through the interstellar medium (ISM), they are subject to various propagation effects, including interstellar scintillation (ISS; see e.g., L. C. Lee & J. R. Jokipii 1975; J. W. Armstrong 1984).
@@
This phenomenon has been extensively studied in compact radio sources such as pulsars and active galactic nuclei (AGN; B. J. Rickett, 1990; J. Goodman, 1997), and has also been observed in several GRB afterglows, (D. A. Frail et al., 1997, 2000; P. Chandra et al., 2008; J. Granot, 2008). ISS arises from small-scale electron density fluctuations in the ISM, which may induce rapid variations in signal intensity if the emitting source is sufficiently compact. Therefore, ISS provide important constraints on the angular size of the emitting region and, consequently, on the Lorentz factor of the outflow (D. S. Heeschen & B. J. Rickett, 1987; J. Dennett-Thorpe & A. G. de Bruyn, 2002; J. E. J. Lovell et al., 2008; B. J. Rickett et al., 1995), helping to determine whether the emission arises from a relativistic outflow.
This paper presents a study of EP241021a, a FXRT discovered by the Wide-field X-ray Telescope (WXT) onboard the EP on October 10, 2024 at $T_{0}=\text{05:07:56 UT}$, with a duration of approximately $100\,\mathrm{s}$ (J. W. Hu et al., 2024b). Notably, no gamma-ray counterpart was detected (D. Svinkin et al., 2024a) and the event was localized at a relatively low redshift of $z=0.75$ (see Fig. 1; G. Pugliese et al., 2024; W. Zheng et al., 2024; I. Pérez-Fournon et al., 2024). We carried out extensive radio follow-up of the afterglow, which remained detectable for over 100 days post-burst in the observer’s frame, placing it among the longest-lived radio afterglows to date (<< 4% of the events; P. Chandra et al., 2008).
@@ 2 Prompt Emission
Our dataset comprises 84 EP events observed up to March 21, 2025. Of these, 55 events have complete prompt emission coverage from the Fermi Gamma-ray Burst Monitor (GBM; C. Meegan et al., 2009) or were detections from Konus-Wind (R. L. Aptekar et al., 1995).
@@
Fig. 2 compares the (observer frame) gamma-ray (10–1000 keV) and X-ray (0.5–4 keV) fluxes for a sample of FXRTs, including EP241021a. For this study, $3\sigma$ upper limits on gamma-ray fluxes were derived using the Gamma-ray Targeted Search (GTS) [^1] https://github.com/USRA-STI/gamma-ray-targeted-search, built around the Gamma-ray Data Tools (A. Goldstein et al., 2024). We adopt a Band function [^2] The Band function is a phenomenological model of GRB spectra (D. Band et al., 1993). See Equation 1. defined as: $N(E)=\begin{cases}A\left(\dfrac{E}{100\,\mathrm{keV}}\right)^{\alpha}\exp\left(-\dfrac{E}{E_{0}}\right),&E<(\alpha-\beta)E_{0}\\ A\left(\dfrac{(\alpha-\beta)E_{0}}{100\,\mathrm{keV}}\right)^{\alpha-\beta}\exp(\beta-\alpha)\left(\dfrac{E}{100\,\mathrm{keV}}\right)^{\beta},&E\geq(\alpha-\beta)E_{0}\end{cases}$ (1) with the peak energy given by $E_{\rm peak}=(2+\alpha)E_{0}$, and we adopt typical GRB parameters $\alpha=-1.0$, $\beta=-2.3$, and $E_{\rm peak}=230~\mathrm{keV}$ (Normal template) to calculate the upper limits of energy fluxes in the case of non-detections. For events with detections, the time-averaged gamma-ray fluxes were taken directly from the values reported through GCN circulars (see Appendix A). The dashed lines in Fig. 2 indicate constant photon indices ($\Gamma=0.5$, $1.0$, and $1.5$), assuming a simple power-law spectrum. We observe that most gamma-ray bright FXRTs lie at $\Gamma\sim 1.0$, consistent with the region typically occupied by standard GRBs (S. Poolakkil et al., 2021).
For EP241021a, we used the X-ray flux (0.5–4 keV) estimated by J. W. Hu et al. (2024b) of $(3.3^{+1.5}_{-1.7})\times 10^{-10}~\mathrm{erg~cm^{-2}~s^{-1}}$. Using the targeted search, we place a $3\sigma$ upper limit on the gamma-ray flux of EP241021a as $3.2\times 10^{-8}~\mathrm{erg~cm^{-2}~s^{-1}}$ for a timescale of 8.192 s. At a redshift of $z=0.75$, this corresponds to a limit on the isotropic-equivalent gamma-ray energy of $E_{\gamma,\rm iso}\lesssim 7.2\times 10^{50}$ erg, consistent with the value inferred by M. Busmann et al. (2025) and at the lower end of the distribution for cosmological GRBs, although still consistent with it.
@@ 3.1 Australia Telescope Compact Array
Australia Telescope Compact Array (ATCA) observations were conducted over 11 epochs between October 28, 2024, and January 25, 2025, covering frequencies from 2.1 GHz to 16.7 GHz. The bandwidth in all the observations was 2 GHz in the Compact Array Broadband Backend 1 MHz per frequency channel mode. The primary and bandpass calibrator was 1934-638 for all observations at 2.1, 5.5 and 9 GHz. At 16.7 GHz, 1921–293 served as the bandpass calibrator observations, while 1934-638 remained the primary calibrator. The phase calibrator 0146+056 was used to correct complex gains. The 21.2 GHz data were taken at the same time of the 16.7 GHz ones but the 21.2 GHz flux densities are not presented in this work because high system temperature values and large flux density scale corrections from the primary calibrator resulted in unreliable flux density values for the phase calibrator and the target source. Data reduction was performed using the Miriad (R. J. Sault et al., 1995) software package, which was used to flag Radio Frequency Interferences (RFIs), calibrate the complex gains, bootstrap the flux density scale, and image the fringe visibilities. A null robustness parameter value was used as a good trade-off between uniform and natural weighting in the imaging strategy. Final cleaned radio maps were inspected by Karma kvis (KARMA kvis, 1995) and ds9 (Smithsonian Astrophysical Observatory, 2000) visualisation softwares. When a detection was found the peak flux was extracted, otherwise, a 3-$\sigma$ upper limit was derived from the Root Mean Square (RMS) noise in the restored maps. The uncertainties on the flux densities were estimated by summing in quadrature the map RMS noise and 5% of the peak flux density, accounting for residual complex gain calibration errors.
@@
EP241021a was observed with enhanced Multi Element Remotely Linked Interferometer Network (e-MERLIN) in four epochs at the centre frequency of 5 GHz with a total bandwidth of 512 MHz divided in 4 sub-bands of 128 MHz each. The data were correlated by the e-MERLIN software correlator. The resulting fringe visibilities were calibrated in the CASA software v.5.8 using the e-MERLIN CASA pipeline (e-MERLIN pipeline, 2019; J. Moldon, 2021) and imaged in CASA v.5.5.0 using a Briggs weighting scheme and a robustness value of 0.5. The target was observed together with 1331+3030 (alias 3C286), 1407+2827, 0154+0823 as the flux calibrator , the pointing calibrator and the bandpass calibrator, respectively. The target was fit with a 2D Gaussian in CASA and the peak flux was quoted in the flux density measurement, being it point-like in the $100\times 40$ mas final cleaned maps. The e-MERLIN position obtained by 2D Gaussian fitting to the cleaned map of EP241021a is RA = 01h 55m 23.4325s, Dec = +05 56 17.774′′ (J2000) with a positional $1\sigma$ uncertainity of (3, 9) mas in RA and DEC, respectively. Flux uncertainties were estimated as detailed in Table 1. Detections and the earlier flux upper limit obtained by this work are shown in Table 2.
3.3 Very Large Array
[…]